Circular No. 8021

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002iu–2002je R. C. Smith and N. Suntzeff, Cerro Tololo Interamerican Observa-

tory (CTIO); and C. Stubbs, University of Washington, on behalf of the ESSENCE project (involving also C. Aguilera, B. Barris, A. Becker, P. Challis, R. Chornock, A. Clocchiatti, R. Covarrubias, A. V. Filippenko, P. Garnavich, S. T. Holland, S. Jha, R. Kirshner, K. Krisciunas, B. Leibundgut, W. Li, T. Matheson, A. Miceli, G. Miknaitis, A. Rest, A. Riess, B. Schmidt, J. Sollerman, J. Spyromilio, and J. Tonry), report the discovery of ten supernovae, each present in at least two bands (V, R, I) taken with the CTIO Blanco 4-m telescope (+ 8K Mosaic II imager) by project members and S. Nikolaev. The supernovae were identified via subtraction of images taken on Oct. 8–14 UT (when none of the new objects was visible, to limiting mag R ∼ 23.5) from those taken on the tabulated discovery dates, with further confirming images taken on Nov. 11 and 13, and with spectroscopy obtained on Nov. 1–11 by project members at the Keck I (+ LRIS), Keck II (+ ESI), and MMT telescopes and by M. Hamuy at the Magellan telescope. SNe 2002iu (z = 0.1), 2002iv (0.2), 2002iy (0.5), 2002iz (0.4), 2002ja (0.3), 2002jb (0.2), 2002jc (0.5), and 2002jd (0.3) are of type Ia; SNe 2002iw (z = 0.2) and 2002ix (0.3) are of type II. Offsets are from the host-galaxy nuclei; in some cases no host was detected. Additional supernova suspects, finding charts, and FITS images are available

at http://www.ctio.noao.edu/essence.

Further to IAUC 8017, B. Beutler and W. Li report the KAIT discovery of SN 2002je. Additional unfiltered magnitudes: Oct. 31.2, [19.0: (KAIT); Nov. 21.1, ∼ 17.5 (M. Schwartz, 0.8-m Tenagra II automated telescope).

SN  2002 UT  α_{2000}  δ_{2000}  R  Offset
2002iu Oct. 30  0^{h}13^{m}33.^{s}10 −10^{o}13′09′′.9 18.9 5′′.6 W, 1′′.4 S
2002iv Oct. 30  2 19 16.11 − 7 44 06.7 20.9 1′′.0 E, 0′′.6 N
2002iw Nov. 1  23 43 07.25 − 9 48 05.2 21.8 2′′.4 W, 0′′.1 N
2002ix Nov. 1  23 31 00.85 − 9 29 07.2 22.2  —
2002iy Nov. 5  2 30 40.00 − 8 11 40.5 21.3  —
2002iz Nov. 5  2 31 20.73 − 8 36 13.1 22.1 1′′.4 W, 0′′.1 N
2002ja Nov. 9  23 30 09.67 − 9 35 01.8 22.2 1′′.0 E, 1′′.3 N
2002jb Nov. 9  23 29 44.14 − 9 36 34.3 21.4 1′′.0 W, 0′′.5 S
2002jc Nov. 9  2 07 27.28 − 3 50 20.7 23.1  —
2002jd Nov. 11  0 28 38.40 + 0 40 29.3 22.0  —
2002je Nov. 20.2  0 27 22.42 +24 10 59.1 17.7 2′′.2 W, 3′′.4 N

2002 November 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green